OpenStudy (anonymous):

The line y = 0.15x + 0.79 represents an estimate of the average cost of gasoline each year. The line 0.11x - y = -0.85 estimates the price of gasoline in January of each year (“Consumer price index,” 2006). a. Do you expect the lines to be intersecting, parallel, or perpendicular? Explain your reasoning.

hero (hero):

You're still doing math?

hero (hero):

What do you think the answer is?

OpenStudy (anonymous):

Unfortunately,my baby has been not letting me finish, now im in danger of failing this assigment

OpenStudy (anonymous):

I dont know. I am never given the time to learn

hero (hero):

How many questions do you have left?

OpenStudy (anonymous):

just a couple

OpenStudy (anonymous):

my baby wont let me finish. I just got her to fall asleep, but who knows for how long

hero (hero):

Oh okay.... Well, to answer the question, the lines will intersect....and the reason is because they both have different slopes. That means that they are definitely not parallel. Also, they are not perpendicular because it order for that to happen their products would have to be equal to -1. Instead, their slopes are very close to each other....both slopes are positive so, if you graph both lines, you will see them intersect at a certain point in the 1st quadrant. You can even go as far as to see the exact point at which they intersect by setting the equations for both lines equal to each other and solving for x, then y as follows: .15x + .79 = .11x + .85 .04x = .10 x = 2.5 To find y, plug x into one of the equations: .15(2.5) +.79 = 1.165 So the point of intersection is (2.5, 1.165)

OpenStudy (anonymous):

b. Use the equations of the lines to determine if they are parallel. What did you find? c. Did your answer to Part b. confirm your expectation in Part a?

hero (hero):

We just determined that they were not parallel. They have different slopes, therefore they are not parallel. Only lines with the same slope have parallel lines. Yes, because we expected the lines to be intersecting and not parallel.

OpenStudy (anonymous):

Can you please answer one more for me?

OpenStudy (anonymous):

Examine the rise in gasoline prices from 1997 to 2006. The price of regular unleaded gasoline in January 1997 was $1.26, and in January 2006, the price of regular unleaded gasoline was $2.31 (“Consumer price index,” 2006). Use the coordinates (1997, 1.26) and (2006, 2.31) to find the slope, or rate of change, between the two points. Describe how you arrived at your answer.

OpenStudy (anonymous):

HERO!

hero (hero):

Oh

hero (hero):

Is this the last one?

OpenStudy (anonymous):

Yes,and Im so sorry

hero (hero):

\[\frac {y_2 - y_1} {x_2 - y_1} = \frac {2.31 - 1.26} {2006 - 1997} = \frac {1.05} {10} = .105\]

hero (hero):

You arrive at the answer by using the formula for slope and inserting the given points into the formula then evaluating